OpenStudy (anonymous):

Solve negative two-thirds times a plus two-fifths equals three-fourths for a

OpenStudy (adamk):

-2a/3 + 2/5 = 3/4?

OpenStudy (adamk):

-2a/3 = 7/20 -2a = 21/20 a = -21/40

OpenStudy (anonymous):

Thank you so much